A local salesman receives a base salary of $950 monthly. He also receives a commission of 8% on all sales over $1750. How much would he have to sell in a month if he needed to have a monthly income of $3000?
Answer by richwmiller(17219)

3000=950+.08*(s-1750)
3000=0.08*(s+10125.)
3000=0.08*s+810.
2190=0.08*s
s =27375